# An Empirical Study on LLM-based Agents for Automated Bug Fixing

### v1

@@ 7.2. Fault Localization
Fault localization (FL) (Wong et al., 2016) techniques aim to discover and analyze the location and causes of faults, which can be categorized into dynamic and static approaches. Dynamic FL techniques, such as spectrum-based fault localization (SBFL) (Abreu et al., 2007, 2009) and mutation-based fault localization (MBFL) (Papadakis and Le Traon, 2015), analyze the dynamic execution information of a program to determine fault locations, though they are resource-intensive. Static FL techniques (Mao et al., 2014) determine fault locations through semantic or syntactic analysis at the bug report or source code level, offering fast detection with low resource consumption. Advanced FL techniques, such as multiple fault localization (MFL) and combined dynamic and static methods, have emerged to guide APR tools in finding and fixing more errors (Xiao et al., 2021; Kim et al., 2019; Neelofar et al., 2017).
7.3. Automated Program Repair
@@
Automated program repair (APR) (Le Goues et al., 2019) has attracted significant attention over the past decade. APR techniques aim to generate patches for buggy programs to pass given test suites. These techniques can be categorized into search-based (Li et al., 2022b; Mehne et al., 2018), semantics-based (Le et al., 2017; Nguyen et al., 2013; Le et al., 2016), and pattern/learning-based approaches (Li et al., 2020, 2022a; Zhang et al., 2023a). Search-based APR techniques like GenProg (Le Goues et al., 2011) use predefined code mutation operators to generate patches, while semantics-based APR techniques generate patches by solving repair constraints based on test suite specifications. Learning-based APR techniques, such as those utilizing deep learning models, train on large code repositories to predict correct patches. Recent work has shown the use of LLMs for APR, often focusing on constructing APR-specific prompts to guide LLMs in generating patches for buggy program statements (Xia et al., 2023).
@@ 8. Conclusion
In this paper, we analyzed the top 4 commercial systems and the top 3 open-source systems on SWE-bench Lite. We conducted detailed analyses of the performance of LLM-based Agents in automatic bug fixing for code repositories, the performance of different systems in Fault Localization, and their performance in Reproduction. The analysis results indicate that to further enhance the capabilities of LLM-based Agents in bug fixing, future efforts should focus on improving the reasoning ability of LLMs. Additionally, attention should be given to the Agentic flow design, considering the quality of issues, stack traces, and the correctness of reproductions.
@@
All raw data comes from the official SWE-bench experiment repository: https://github.com/swe-bench/experiments/

### v2

@@ 7.2. Fault Localization and Program Repair
Fault localization (FL) (Wong et al., 2016) techniques aim to discover and analyze the location and causes of faults, which can be categorized into dynamic and static approaches. Dynamic techniques, such as spectrum-based fault localization (SBFL) (Abreu et al., 2007, 2009) and mutation-based fault localization (MBFL) (Papadakis and Le Traon, 2015), analyze the dynamic execution information to determine fault locations, though they are resource-intensive. Static FL techniques (Mao et al., 2014) determine fault locations through semantic or syntactic analysis at the bug report or source code level, offering fast detection with low resource consumption. Advanced FL techniques, such as multiple fault localization (MFL) and combined dynamic and static methods, have emerged to find more errors (Xiao et al., 2021; Kim et al., 2019; Neelofar et al., 2017).
@@
Automated program repair (APR) (Le Goues et al., 2019) has attracted significant attention over the past decade. APR techniques aim to generate patches for buggy programs to pass given test suites. These techniques can be categorized into search-based (Li et al., 2022b; Mehne et al., 2018), semantics-based (Le et al., 2017; Nguyen et al., 2013; Le et al., 2016), and pattern/learning-based approaches (Li et al., 2020, 2022a; Zhang et al., 2023). Search-based APR techniques like GenProg (Le Goues et al., 2011) use predefined code mutation operators to generate patches, while semantics-based APR techniques generate patches by solving repair constraints based on test suite specifications. Learning-based APR techniques, such as those utilizing deep learning models, train on large code repositories to predict correct patches. Recent work uses LLMs for APR, often focusing on constructing APR-specific prompts to guide LLMs in generating patches for buggy program statements (Xia et al., 2023).
@@ 8. Conclusion
In this paper, we analyzed the top 6 bug-fixing systems on SWE-bench Verified. We conducted detailed analyses of the performance of LLM-based Agents in automatic bug fixing for code repositories, and the performance of different systems in different levels of fault localization. Moreover, we developed a mainstream workflow bug-fixing agent to evaluate the overall performance on bug reproduction tasks. The analysis results indicate that to further enhance the capabilities of LLM-based Agents in bug fixing, future efforts should focus on improving the reasoning ability of LLMs, and the Agentic flow design, considering the quality of issues, stack traces, and the correctness of reproductions.
@@
The code and related experimental data in this paper are accessible on this page: https://github.com/ResearchOpenRepos/bug_fixing_agent_empirical_study